When released from rest in a uniform electric field of magnitude 1.00 x 10^4 N/C, a charged particle travels 2.00 cm in 2.00 x 104-7s in the direction of the field. You can ignore any nonelectric forces acting on the particle. Set up your solution to answer the following questions: (0 Is the particle positively or negatively charged? (ii) What is the charge-to-mass ratio of this particle, i.e. the ratio of the charge to the mass? (i) Other experiments show that the mass of this particle 6.64 x 10 -27 kg. What is the charge of the particle? Apply the following steps in your solution: a) Describe the approach you will take b) Visualize the problem c) Set up the equations you will need to solve the problem d) Apply algebraic techniques to şolve for the unknown of interest e) Calculate the answers to parts (i) -(ii)
Answer #1

(i)

The particle is accelerating in the direction of electric field.

so, the particle is positive.

(ii)

electric field apply a electric force on particle.

F = qE

ma = qE

a = qE / m ..........(1)

From kinematic equation,

Distance traveled by the particle, s = ut + (1/2)at^2

since, it starts from rest, u = 0

so s = (1/2)at^2

Put the value of a from eq (1),

s = (1/2)*(qE / m)*t^2

Put the given values,

0.02 = (1/2)*(q / m)*1*10^4*(2*10-7)^2

q / m = 1*10^8 C/kg

(iii)

as given that, m = 6.64*10^(-27) kg

q / 6.64*10^(-27) = 1*10^8

q = 6.64*10^(-19) C